3. Determine the reactions at the supports and then draw a FBD at joint A, B, and 400 N
Answer #1

Solu Hon given that DI POD MOON → FAB 2450 FBD at Joint FAE -Dx =0 Dx=0] to &ty=0 Ay - 2000 Ay+ by = youn Ay = 400-200 (@=0Solution 1 pFy =0 Ay tay = 600N (ONA =0 - 600 (2) + Cy Ca)=0 Cycas = 1200 kom y = 12000- ung (g = 300w FBD at Joint A Ay tay